solve
3x-y=-2
4y=12x-16

Answer by unlockmath(1688)

Hello,
Let's rewrite this as:
3x-y=-2
-12x+4y=-16
Now Multiply the first by 4 to get:
12x-4y=-8
-12x+4y=-16
Add these together to get:
0=-24
Obviously this makes no sense, so there's no solution.
Got it?
RJ